# GCRPNet: Graph-Enhanced Contextual and Regional Perception Network for Salient Object Detection in Optical Remote Sensing Images

### v3

@@ III METHODOLOGY
In this section, we propose a novel architecture named GCRPNet, whose overall structure is illustrated in Fig.2. GCRPNet adopts a typical encoder-decoder framework and integrates three innovative modules: DS-HGAM, MGAM, and LESS2D. To clearly present our work, we first provide a summary and review of existing SSM methods in Section III-A. Then, we introduce the overall architecture of GCRPNet, followed by detailed descriptions of the three core modules, including their design principles and implementation details. Finally, we explain the loss functions used in our approach.
III-A Preliminaries
@@ III-B Overall Structure
The overall structure of GCRPNet, illustrated in Fig.2 (A), the overall architecture of GCRPNet adopts a U-shaped design and consists of three main components: (1) A four-stage VSS encoder built on the VMamba backbone, responsible for multi-scale feature extraction; (2) Four difference-similarity guided hierarchical graph attention module (DS-HGAM) embedded in the skip connections, which fuse the multi-scale features from the encoder and apply graph attention mechanisms to model them; (3) A four-stage locally enhanced VSS (LEVSS) decoder, designed to enhance local modeling and progressively decode the features to obtain fine-grained segmentation maps. In the encoder stage, the input image $I\in\mathbb{R}^{H\times W\times 3}$ first passes through a patch embedding module composed of two convolutional layers, which performs patch-wise embedding. Each encoder stage $s\in\{1,2,3,4\}$ contains a $2\times$ downsampling layer followed by multiple stacked VSS blocks for extracting multi-scale features $F_{i}$ where $i\in\{1,2,3,4\}$. After the encoder, the features from the four stages are fed into the difference-similarity guided hierarchical graph attention module (DS-HGAM) in a cascaded manner. These modules process multi-scale features and integrate them into their respective feature representations. Through the graph attention mechanism, the feature $F^{g}_{i}$ is obtained, which captures long-range relationships among pixels in the fused feature maps, thereby enabling the model to identify salient objects in the presence of severe background interference and enhance its discriminative capability for salient regions. In the decoder, the network comprises four decoding stages, each stage $D_{s}$, where $s\in\{4,3,2,1\}$ incorporates multiple stacked LEVSS blocks to enhance local feature modeling capability. The first decoding stage $D_{4}$ directly takes the skip-connected features $F^{g}_{4}$ from the encoder as input. The remaining three stages $s\in\{3,2,1\}$ concatenate high-level skip-connected features $F^{g}_{s}$ with the upsampled intermediate features from the previous stage along the channel dimension to form the fused input features.
III-C Difference-Similarity Guided Hierarchical Graph Attention Module
Traditional skip connections in U-shaped network architectures, due to their simplistic and direct feature concatenation strategy, which lack modeling of spatial context and fine-grained details, often fail to effectively distinguish foreground from background in low-contrast scenarios, leading to blurred boundaries and inaccurate localization. Notably, although ORSIs suffer from complex backgrounds and low contrast, salient objects often exhibit a clustered distribution with similar visual patterns across regions. To better exploit this property, we propose a DS-HGAM. Instead of directly concatenating encoder and decoder features, DS-HGAM performs multi-scale fusion of skip-connected features and employs a graph neural network to model spatial relationships and contextual dependencies across hierarchical levels. By computing attention across regions and scales, DS-HGAM effectively captures complementary information among visually similar patches at different resolutions. This not only enhances the model’s structural perception capability but also strengthens the representation of fine-grained details, enabling more accurate saliency prediction in low-contrast and cluttered scenes.
Specifically, DS-HGAM(as illustrated in Fig.3) first targets the four groups of skip-connected features $F_{i}$, where $i\in\{1,2,3,4\}$ from different encoding layers and introduces multi-scale feature fusion (MSFF) modules for each. These modules perform spatial alignment on the multi-scale features, fusing them into features of corresponding resolutions. The feature fusion process is as follows:
$F^{\prime}_{i}=Conv_{1\times 1}(Concat(\phi_{1}(F_{1}),\phi_{2}(F_{2}),\phi_{3}(F_{3}),\phi_{4}(F_{4}))$ (7)
@@ III-C Difference-Similarity Guided Hierarchical Graph Attention Module
Where $\phi_{i}(\cdot)$ denotes the operation of adjusting features to the resolution of the $i$ layer via upsampling or downsampling; $Concat$ represents channel-wise concatenation; $Conv_{1\times 1}$ denotes a convolution operation used to compress the number of channels after concatenation.
Figure 3: Illustration of DS-HGAM
Subsequently, the RGCA module takes the fused feature $F^{\prime}_{i}$ as input, where $i\in\{1,2,3,4\}$. Feature $F^{\prime}_{i}$ is first fed into the CCS module. The CCS module combines channel attention and spatial attention to enhance the focus on salient regions within the target, followed by convolution operations to process the fused features and enhance their representational power.
$F^{c}_{i}=CCS(F^{\prime}_{i})$ (8)
@@
Meanwhile, to reduce the number of parameters processed by the subsequent graph attenton networks operations, the RGCA module downsamples the fused feature, the feature $F^{c}_{i}\in\mathbb{R}^{h\times w\times c}$ is fed into a convolutional layer with a kernel size of $s\times s$ and a stride of $s$. Then, each row of pixels in the output feature $F^{c^{\prime}}_{i}\in\mathbb{R}^{h/s\times w/s\times c}$ is treated as a graph node $f_{i}\in\mathbb{R}^{d}$. The graph composed of image patches can be represented as: $F_{in}=\{\vec{f}_{1},\vec{f}_{2},...,\vec{f}_{n}\},n=h/s\times w/s$. To transform the input features into higher-level representations and enhance the expressiveness of the network, we apply a shared learnable weight matrix $W\in\mathbb{R}^{d^{\prime}\times d}$ to linearly transform each node. Then, a shared self-attention weight matrix $l\in\mathbb{R}^{2d^{\prime}\times 1}$ is used to compute the attention coefficients between node pairs. We only compute $a_{ij}$ for node $j$, where $\mathcal{N}_{i}$ denotes the set of first-order neighboring nodes of node $i$. To better compare the attention coefficients $a$ between different nodes, we apply the softmax function to normalize all the selected $j$. Therefore, the formula for $a_{ij}$ is:
@@
The final output feature has the same dimensionality as the input feature, denoted as $F_{out}=\{\vec{f}^{\prime}_{1},\vec{f}^{\prime}_{2},...,\vec{f}^{\prime}_{n}\},\vec{f}^{\prime}_{n}\in\mathbb{R}^{d}$. This output feature $F_{out}$ is then passed through a deconvolution and upsampling operation, and multiplied with the obtained channel attention weight $w$ to form the residual information flow of $F_{i}$, thereby compensating for the information lost during downsampling.
The overall operation process is as follows:
$F^{gat}_{i}=GAT(F^{c}_{i})\otimes CA(F^{c}_{i})$ (11)
$F^{g}_{i}=F^{gat}_{i}+F_{i}$ (12)
Finally, fuse the original feature $F_{i}$ with the enhanced feature $F^{gat}_{i}$ output through a residual connection to produce the refined feature representation $F^{g}_{i}$.
III-D Locally Enhanced Visual State Space
Salient objects in ORSIs often exhibit significant variations in scale and complex structural characteristics. This requires the model to possess strong capabilities in both local region and multi-scale representation in order to achieve accurate saliency detection. However, the traditional Mamba scanning strategy flattens 2D images into 1D sequences, disrupting local region spatial dependencies and making it difficult for the model to capture fine details. How to enable the Mamba model to effectively model local region is still a challenge. Therefore, we propose the LEVSS module, which is designed on top of the original Mamba module to enhance the representation of local region information and significantly improve the model’s ability to capture local region dependencies. The structure of LEVSS is illustrated in Fig.2 (B) and consists of two main components: MCAEM and LESS2D.
III-D1 Multi-Granularity Collaborative Attention Enhancement Module
In the ORSI-SOD task, effectively adapting to multi-scale object variations is essential for robust saliency detection. To this end, we propose a MCAEM that effectively combines multi-scale feature extraction with channel and spatial attention mechanisms to enhance feature representation (as illustrated in Fig.4). This module is integrated into the LEVSS block and applied before sequence scanning, aiming to reinforce multi-scale feature representation and compensate for the limited local region perception capability of sequence modeling structures.
Specifically, MCAEM employs depthwise separable convolutions[53] with three different kernel sizes $(3\times 3,5\times 5,7\times 7)$ to extract features at multiple receptive fields, thereby strengthening the representation of structural information across scales. In parallel, channel attention and spatial attention mechanisms are applied to the extracted features for adaptive enhancement. Channel attention adjusts the responses of different semantic types ($e.g.$, edges, textures, colors) from a global perspective, which helps improve detection accuracy. Meanwhile, spatial attention focuses on localizing important regions in the image and emphasizes discriminative areas, while the parallel design helps mitigate redundancy and reduce biased dependencies between feature channels or spatial locations.
Figure 4: Illustration of the proposed MCAEM
Given an input feature map $F_{i}\in\mathbb{R}^{h\times w\times c}$, the overall processing of MCAEM can be summarized in the following steps:
$F^{\prime}_{k}=DWConv_{k\times k}(F),k\in\{3,5,7\}$ (13)
$F^{att}_{k}=CA(F^{\prime}_{k})+SA(F^{\prime}_{k})$ (14)
$F_{out}=F+Conv_{1\times 1}(Concat(F^{att}_{3},F^{att}_{5},F^{att}_{7}))$ (15)
Here, $CA(\cdot)$ and $SA(\cdot)$ refer to the channel and spatial attention modules respectively, $Conv_{1\times 1}$ represents a $1\times 1$ convolution, and $Concat$ indicates concatenation along the channel dimension.
III-D2 Locally Enhanced 2D Selective Scan
In salient object detection of complex objects, accurately modeling local details plays a vital role in improving detection accuracy. To this end, we propose a LESS2D module, which aims to address the limitations of traditional SS2D in capturing local features, while simultaneously balancing the Mamba model’s global context modeling capability with enhanced sensitivity to local details.
Specifically, the core motivation of this module lies in the observation that as the resolution of the feature map increases, its spatial geometric structure becomes more complex, and the contained local details grow richer. Flattening a 2D image into a 1D sequence not only disrupts the original spatial adjacency but also causes local region features to become dispersed in the sequence, thereby weakening the model’s ability to capture local dependencies. To address this issue, we propose LESS2D (as illustrated in Fig.2 (B)) and integrate it into a four-stage decoder architecture. Given an input image $I\in\mathbb{R}^{h\times w\times c}$, we perform the split patch operation on the extracted multi-scale feature maps, dividing them into multiple image blocks to facilitate subsequent local feature modeling and processing. In our method, for low-resolution feature maps ($e.g.$, $\frac{1}{16}h\times\frac{1}{16}w$), due to their limited spatial modeling range and concentrated geometric relationships, LESS2D applies a global four-directional scanning strategy to fully retain long-range dependency modeling capability. In contrast, for medium and high resolution feature maps $(e.g.,\frac{1}{8}h\times\frac{1}{8}w,\frac{1}{4}h\times\frac{1}{4}w,\frac{1}{2}h\times\frac{1}{2}w)$, the image is divided into multiple non-overlapping local patches (e.g., $4$, 1616, and $64$ patches, respectively), and four-directional scanning is performed independently within each patch. The scanning sequences from all patches are then concatenated in order to form a complete sequence, which is subsequently fed into downstream modules for further processing.
Taking a feature map $F$ of size $\mathbf{\frac{1}{4}h\times\frac{1}{4}w}$ as an example, we divide it into 16 non-overlapping sub-regions, each denoted as $F_{i}$, where $i\in\{1,2,...,16\}$. The overall process is carried out as follows:
@@ III-D2 Locally Enhanced 2D Selective Scan
$\vec{S}=Concat(Scan^{\rightarrow}(F_{1}),Scan^{\rightarrow}(F_{2}),...,Scan^{\rightarrow}(F_{16}))$ (17)
Each block contributes four directional scans denoted as $^{\rightarrow}, ScanScan^{\downarrow}$, $Scan^{\leftarrow}$, and $Scan^{\uparrow}$ (as illustrated in Fig.2 (C)), which represent left-to-right, top-to-bottom, right-to-left, and bottom-to-top scanning, respectively. Eq. 17 illustrates the left-to-right scanning process, where $\vec{S}$ represents the complete image sequence.
III-E Loss Function

### v4

@@ 3 Methodology
In this section, we propose a novel architecture named ASCNet, whose overall structure is illustrated in Figure 3. ASCNet adopts a typical encoder-decoder framework and integrates three innovative modules: MLCM, GPM, and DAGS. To clearly present our work, we first provide a summary and review of existing SSM methods in Section III-A. Then, we introduce the overall architecture of ASCNet, followed by detailed descriptions of the three core modules, including their design principles and implementation details. Finally, we explain the loss functions used in our approach.
3.1 Preliminaries
@@ 3.2 Overall Structure
The overall architecture of ASCNet is illustrated in Figure 3(A). ASCNet is composed of three main components: (1) a four-stage VSS encoder for hierarchical multi-scale feature extraction; (2) four Multi-Level Context Modules (MLCMs) that integrate features across different scales and model contextual relationships via topology-aware attention mechanisms; and (3) a four-stage APVSS decoder, which enhances local feature modeling and progressively restores spatial resolution to generate fine-grained segmentation results. In the encoding stage, the input image $I\in\mathbb{R}^{H\times W\times 3}$ is first fed into a patch embedding module consisting of two convolutional layers, which transforms the image into patch-wise feature representations. The encoder is organized into four stages $s\in\{1,2,3,4\}$, where each stage contains a $2\times$ downsampling operation followed by multiple stacked VSS blocks (the number of blocks in each stage are respectively 2, 2, 15, 2). Through this hierarchical design, multi-scale feature maps $F_{i},\ i\in\{1,2,3,4\}$ are extracted at different spatial resolutions. After encoding, the multi-scale features are successively processed by the MLCMs in a cascaded manner. Each MLCM aligns and fuses features from different levels and employs a topology-aware attention mechanism to explicitly model long-range dependencies among regions in the fused feature maps. As a result, the enhanced feature representations $F_{i}^{\prime}$ are obtained, which effectively capture global contextual relationships. This design enables the network to reliably highlight salient objects under complex background interference and significantly improves the discriminative capability for salient regions. In the decoding stage, ASCNet consists of four decoding levels. Each decoding stage contains two stacked Adaptive Patchwise Visual State Spac (APVSS) blocks to further strengthen local feature modeling within each scale. The deepest decoding stage takes the high-level encoded features as input, while the remaining stages integrate the context-enhanced features at the corresponding scale with the upsampled intermediate representations from the previous decoding stage along the channel dimension. Through progressive feature fusion and refinement, the decoder gradually recovers spatial details and produces high-resolution segmentation predictions.
3.3 Multi-Level Context Module
In ORSI-SOD, representations from a single abstraction level are often insufficient for reliable saliency discrimination in complex and low-contrast scenes. High-level features provide coarse semantic cues, while low-level features preserve fine-grained spatial structures and boundary details, making multi-level contextual aggregation essential for accurate salient object localization. Moreover, although optical ORSIs often exhibit visually cluttered backgrounds, salient objects tend to present spatially clustered distributions with consistent visual patterns across regions and representation levels. This property suggests that saliency discrimination can benefit from explicitly modeling relationships among spatially separated yet semantically related regions across multiple levels. Motivated by this observation, we introduce the Multi-Level Context Module (MLCM), which aims to aggregate contextual cues across different representation levels and capture region-level dependencies through structured relational modeling.
As illustrated in Figure 3(D), MLCM operates on a set of feature representations $F_{i}$, where $i\in\{1,2,3,4\}$ extracted from different encoding stages, each corresponding to a distinct abstraction level. For each level, a multi-scale feature fusion operation is first employed to align and integrate information from its adjacent representation levels into a unified representation at the target resolution. This design enables each level to access complementary contextual cues from other abstraction levels. The fusion process is formulated as:
$F_{i}^{c}=Conv_{1\times 1}(Concat(\phi_{1}(F_{i-1}),\phi_{3}(F_{i}),\phi_{4}(F_{i+1}))$ (7)
@@ 3.3 Multi-Level Context Module
where $\phi_{i}(\cdot)$ denotes resolution alignment via upsampling or downsampling, $Concat$ represents channel-wise concatenation, and $Conv_{1\times 1}$ compresses the fused representation into a compact feature space.
@@
Meanwhile, to reduce the number of parameters processed by the subsequent topology-aware attention networks operations, the module downsamples the fused feature, the feature $F^{c}_{i}\in\mathbb{R}^{h\times w\times c}$ is fed into a convolutional layer with a kernel size of $s\times s$ and a stride of $s$ ($s=3$). Then, each spatial location in the downsampled feature map $F^{c^{\prime}}_{i}\in\mathbb{R}^{h/s\times w/s\times c}$ (corresponding to an $s\times s$ local region in the original feature) is treated as a topological node, and the channel vector at that location is used as the node feature $\vec{f}_{k}\in\mathbb{R}^{d}$. Therefore, the downsampled feature map can be rearranged into a region-level feature representation with explicit spatial topological structure: $F_{in}=\{\vec{f}_{1},\vec{f}_{2},...,\vec{f}_{n}\},n=h/s\times w/s$. To transform the input features into higher-level representations and enhance the expressiveness of the network, we apply a shared learnable weight matrix $W\in\mathbb{R}^{d^{\prime}\times d}$ to linearly transform each node. Then, a shared self-attention weight matrix $l\in\mathbb{R}^{2d^{\prime}\times 1}$ is used to compute the attention coefficients between node pairs. The attention coefficient $a_{ij}$ is computed only for nodes $j\in\mathcal{N}_{i}$, where $\mathcal{N}_{i}$ denotes the first-order neighborhood of node $i$ in the reorganized region-level representation. Specifically, $\mathcal{N}_{i}$ consists of the spatially adjacent nodes in the 2D grid of the downsampled feature map, implemented using a fixed local neighborhood. This localized attention design enables efficient relational modeling while preserving local spatial consistency. To better compare the attention coefficients $a$ between different nodes, we apply the softmax function to normalize all the selected $j$. Therefore, the formula for $a_{ij}$ is:
@@
Finally, the feature maps enhanced by two consecutive topological dependency modeling steps are reshaped back to the spatial domain and projected to the original resolution through a deconvolution followed by an upsampling operation, denoted as $F_{out}=\{\vec{f}^{\prime}_{1},\vec{f}^{\prime}_{2},...,\vec{f}^{\prime}_{n}\},\vec{f}^{\prime}_{n}\in\mathbb{R}^{d}$, yielding the final output of the MLCM. This output feature effectively integrates multi-level contextual information and region-level relational dependencies, serving as a discriminative representation for subsequent saliency decoding.
3.4 Adaptive Patchwise Visual State Space Block
We propose the Adaptive Patchwise Visual State Space (APVSS) module, which is designed on top of the original VMamba module to enhance the representation of local region information and significantly improve the model’s ability to capture local region dependencies. The structure of APVSS is illustrated in Figure 3 (B) and consists of two main components: GPM and DAGS.
3.4.1 Granularity-aware Propagation Module
In optical remote sensing salient object detection, accurate saliency discrimination requires regional representations to be interpreted under globally consistent semantic cues. To this end, we propose the Granularity-aware Propagation Module (GPM), which performs semantic conditioning on regional representations prior to sequence construction, thereby guiding subsequent state space modeling with stable global context (as illustrated in Figure 3(D)). The GPM is embedded into the adaptive state space decoding pipeline and applied before sequence construction. Its role is to establish a semantically conditioned perceptual space in which regional representations can be interpreted more consistently across complex scenes.
Given an input feature map $F_{i}\in\mathbb{R}^{h\times w\times c}$, GPM first partitions it into a set of non-overlapping local patches. Each patch is flattened into a sequence of tokens $T_{i}\in\mathbb{R}^{p^{2}\times C}$, where $p$ ($p=16$) denotes the patch size and $i$ indexes the patch. To inject global semantic priors into regional perception, we introduce 2 learnable global tokens $G\in\mathbb{R}^{2\times C}$ and concatenate them with the patch tokens. The concatenated tokens are then processed by a global semantic interaction operator, which enables feature mixing and information propagation under global semantic conditioning:
$\hat{T}_{i}=[G;T_{i}]\in\mathbb{R}^{(2+p^{2})\times C}$ (10)
$Z_{i}=(LN(\hat{T}_{i})W_{1})\odot\sigma(Pool_{g}(LN(\hat{T}_{i}))W_{2})$ (11)
where $W_{1},W_{2}\in\mathbb{R}^{C\times C}$are learnable projection matrices, $Pool_{g}(\cdot)$denotes global pooling over the token dimension (including global tokens), and $\sigma(\cdot)$is the sigmoid activation function. Through this semantic interaction process, the global tokens serve as semantic anchors that aggregate image-level contextual information and generate global conditioning signals to regulate regional feature responses. As a result, patch tokens are interpreted under globally consistent semantic cues, which enhances the separability between salient objects and background regions in complex spatial layouts. After semantic interaction, only the updated patch tokens are retained:
$T^{\prime}_{i}=Z_{i}[2+1:2+p^{2}]$ (12)
The refined patch tokens are subsequently reshaped back into two-dimensional feature blocks and reassembled to form an enhanced feature map $X^{\prime}$. On top of $X^{\prime}$, a gated convolutional feed-forward network is applied to further regulate feature responses. Specifically, depthwise separable convolution[6] captures local structural patterns, while a gating mechanism adaptively modulates feature activation strength:
$Y=DWConv(X^{\prime})\odot\sigma(X^{\prime})$ (13)
The resulting representation integrates region-level structural information with global semantic conditioning and is forwarded to subsequent sequence construction and state space modeling stages.
3.4.2 Dynamic Adaptive Granularity Scan
Compared with natural images, salient targets in ORSIs are often embedded in cluttered backgrounds and exhibit diverse geometric patterns across scales and regions. Consequently, effective state space modeling for salient object detection in ORSIs requires the sequence construction process to adapt not only to the heterogeneous spatial organization of visual representations across different resolutions, but also to the diverse geometric patterns within local regions. Feature maps at different scales exhibit distinct levels of geometric complexity: lower-resolution representations mainly encode coarse semantic layouts with compact spatial structures, while higher-resolution ones preserve richer geometric details and fine-grained spatial variations [3, 28]. Moreover, even within the same resolution, local regions may present different directional structures and spatial arrangements. However, applying a uniform scanning strategy across all resolutions and spatial locations implicitly assumes homogeneous spatial organization, which is inconsistent with the resolution- and content-dependent nature of dense visual representations. Motivated by this observation, we propose Dynamic Adaptive Granularity Scan (DAGS), a resolution-aware and content-adaptive scanning strategy embedded within the adaptive state space decoder block to better align sequence construction with multi-resolution and region-specific geometric characteristics.
DAGS addresses this issue by dynamically adjusting both the spatial granularity and the relative importance of directional scans during sequence construction. Given an input feature map $I\in\mathbb{R}^{h\times w\times c}$, DAGS first partitions the representation into spatial blocks according to its resolution, enabling resolution-aware modeling of spatial dependencies within the decoder block. For low-resolution feature maps (e.g., $\frac{1}{16}h\times\frac{1}{16}w$), where spatial structures are relatively compact and global semantic layouts dominate, DAGS applies four-directional global scanning to preserve long-range contextual interactions. In contrast, for medium- and high-resolution feature maps (e.g., $\frac{1}{8}h\times\frac{1}{8}w$, $\frac{1}{4}h\times\frac{1}{4}w$, and $\frac{1}{2}h\times\frac{1}{2}w$), the feature maps are divided into multiple non-overlapping sub-regions (e.g., 4, 16, and 64 regions, respectively), allowing localized modeling of fine-grained spatial structures. For each sub-region $F_{i}$, DAGS performs four directional scan operations in parallel:
@@ 3.4.2 Dynamic Adaptive Granularity Scan
Instead of treating all directional scans equally, DAGS introduces a lightweight content-adaptive weighting mechanism to modulate the contribution of each directional sequence. Specifically, a direction-aware response vector $g_{i}\in\mathbb{R}^{4}$ is predicted from the aggregated sub-region feature using global average pooling followed by a learnable projection. The directional importance weights are then obtained via Softmax normalization:
$\mathbf{g}_{i}=\phi\!\left(GAP(F_{i})\right),$ (15)
$w_{i}^{d}=\frac{\exp\left(g_{i}^{d}\right)}{\sum_{d^{\prime}\in\{\rightarrow,\,\downarrow,\,\leftarrow,\,\uparrow\}}\exp\left(g_{i}^{d^{\prime}}\right)},\quad d\in\{\rightarrow,\,\downarrow,\,\leftarrow$$\uparrow\}$. (16)
where $w_{i}^{d}\in\mathbb{R}^{4}$, $\phi(\cdot)$ denotes a learnable projection, and $GAP(\cdot)$ is global average pooling. Each directional scan sequence is then reweighted accordingly:
$\tilde{S}_{i}^{d}=w_{i}^{d}\cdot Scan^{d}(F_{i}),\quad d\in\{\rightarrow,\,\downarrow,\,\leftarrow$$\uparrow\}$. (17)
Finally, the reweighted directional scan sequences from all sub-regions are concatenated in a fixed spatial and directional order to form the complete sequence representation fed into the state space model. By preserving the four-directional scanning paradigm while enabling content-adaptive modulation at the sub-region level, DAGS allows the state space model to flexibly emphasize geometrically consistent directions across different resolutions and spatial locations, effectively balancing global contextual modeling and fine-grained directional sensitivity without introducing additional architectural stages.
3.5 Loss Function